after receiving a 5% pay cut, mario is now earning 10.45 an hour. find is old pay
Answer by Alan3354(69443)

after receiving a 5% pay cut, mario is now earning 10.45 an hour. find is old pay
-----------
10.45/hr = 95% = 0.95*old rate
old rate = 10.45/0.95